# Enhancing the accuracy of observable distributions for galaxies classified in the Projected Phase Space Diagram

### v1

@@ Abstract
Aims. We propose a method to statistically correct the determination of galaxy properties’ distributions accounting for the contamination caused by misclassified galaxies from other classes.
Methods. Using a sample of massive clusters and galaxies in their surroundings taken from the MultiDark Planck 2 simulation combined with the semi-analytic model of galaxy formation sag, we compute the confusion matrix associated to a classification scheme in the PPSD. Based on positions in the PPSD, galaxies are classified as cluster members, backsplash galaxies, recent infallers, infalling galaxies, and interlopers. This classification is determined using probabilities calculated by the code roger  along with a threshold criterion. By inverting the confusion matrix, we are able to get better determinations of distributions of galaxy properties such as colour.
Results. Compared to a direct estimation based solely on the predicted galaxy classes, our method provides better estimates of the mass-dependent colour distribution for the galaxy classes most affected by misclassification: cluster members, backsplash galaxies, and recent infallers. We apply the method to a sample of observed X-ray clusters and galaxies.
Conclusions. Our method can be applied to any classification of galaxies in the PPSD, and to any other galaxy property besides colour, provided an estimation of the confusion matrix. Blue, low-mass galaxies in clusters are almost exclusively recent infaller galaxies that have not yet been quenched by the environmental action of the cluster. Backsplash galaxies are on average redder than expected.
@@ 1 Introduction
In recent years, the use of the Projected Phase Space Diagram (PPSD) has increased as a tool for classifying galaxies according to their kinematic behaviours and dynamical state in and around systems of galaxies such as clusters and groups. The PPSD is a two-dimensional space that combines the projected cluster-centric distance (or group-centric), with the line-of- sight velocity relative to the cluster (or group). Several authors have utilised the PPSD to study the evolution of galaxies. For example, Mahajan et al. (2011) study the decrease in star formation of backsplash galaxies. Muzzin et al. (2014) focus on the quenching of galaxies at $z\sim 1$. Muriel & Coenda (2014) investigate the properties of high and low-velocity galaxies in the outskirts of galaxy clusters. Hernández-Fernández et al. (2014) and Jaffé et al. (2015) study the gas fraction and ram-pressure of galaxies in clusters. Oman & Hudson (2016) study the relationship between the star formation rates of a sample of observed galaxies and their likely orbital histories. Smith et al. (2019) analyse the stellar mass growth histories of galaxies as a function of their infall time. Martínez et al. (2023) study the quenching of galaxies in a sample of X-ray clusters. Aguerri et al. (2023) focus in the properties of barred galaxies with the environment. Sampaio et al. (2024) examine the evolution of blue cloud, green valley, and red-sequence fractions as a function of time since infall. Muñoz Rodríguez et al. (2024) study the impact on black hole growth as a function of the environment.
@@
The aforementioned studies classify galaxies in systems and their outskirts following methods such as those proposed by Rhee et al. (2017), Pasquali et al. (2019), or de los Rios et al. (2021). Based on cosmological hydrodynamic N-body simulations of groups and clusters (Choi & Yi 2017), Rhee et al. (2017) categorised galaxies within PPSD diagram based on the time since their infall into the system. This categorisation included galaxies identified as first, recent, intermediate, and ancient infallers. They delineated specific regions within the PPSD where each of these types of galaxies are preferentially located. Pasquali et al. (2019) present an alternative approach using the same cosmological simulations (Choi & Yi 2017) of groups and clusters employed by Rhee et al. (2017). They establish eight zones of constant mean infall time to investigate environmental influences on satellite galaxies. These zones are defined by analytical curves that characterise the timing of galaxy infall into the system.
@@
Recently, de los Rios et al. (2021) presented the code roger [^1] https://github.com/Martindelosrios/ROGER which employs three different machine learning techniques to classify galaxies based on their projected phase-space positions within and around clusters. They utilised a dataset comprising massive and isolated galaxy clusters from the MultiDark Planck 2 simulation (Klypin et al. 2016). This code establishes a connection between the two-dimensional (2D) phase space position of galaxies and their three-dimensional (3D) orbital classifications. roger  calculates the probabilities of galaxies belonging to any of the five orbital classes: cluster members (CL), backsplash galaxies (BS), recent infallers (RIN), infallers (IN) and interlopers (ITL).
@@
Regardless the chosen method for classifying galaxies in the PPSD diagram, there will inevitably be some degree of contamination among the classes, particularly in regions closer to the centres of clusters. According to Coenda et al. (2022), misclassification have significant implications for interpreting the properties of galaxies in the 2D classes. Therefore, the degree of contamination significantly impacts the conclusions that can be drawn. They emphasise the importance of distinguishing between red and blue galaxy populations to achieve a more precise analysis of observational data. Furthermore, they argue that the 2D analysis reliably provides results only for recent infallers, infalling galaxies, and interloper galaxies within the blue population.
@@
To improve determinations of distributions of galaxy observables such as colours, specific star formation rate, etc., for galaxies that have been classified using their position in the PPSD, in this paper we propose a simple method that reconstruct the intrinsic distribution of galaxy properties. The method relies in an statistic estimation of how well the classification is performed. This paper is organised as follows: in Sect. 2 we present our method; in Sect. 3 we test the method upon a sample of simulated clusters and galaxies; in Sect. 4 we use the method to reconstruct the distributions of colours of galaxies of different classes in and around a sample of X-ray clusters of galaxies; finally, we present our conclusions and mention possible further uses of the method in Sect. 5.
@@ 2 Method
Let us consider $n$ classes of galaxies according to the particular classification scheme we are using over the PPSD. The intrinsic class of each galaxy cannot be uniquely determined by its position in it. At best, we can assign a particular class to a galaxy based on its position in the PPSD using our chosen methodology. In our case, this involves utilising the output probabilities of roger combined with a numerical criterion on their values, as we explain in Sect. 3. This way, each galaxy is assigned a predicted class. Due to uncertainties inherent to the classification procedure, the number of galaxies in the $i-$th predicted class ($i=1,\ldots,n$), $N^{\mathrm{P}}_{i}$, is built up from the contribution of galaxies, in principle, from all intrinsic classes, that were classified as class $i$:
@@
where $N^{\mathrm{I}}_{j}$ is the number of galaxies in our sample that belong to the $j-$th intrinsic class, and $C_{ij}$ is the fraction of galaxies of the $j-$th intrinsic class that were classified as being part of the $i-$th predicted class. The ideal case is $C_{ij}=\delta_{ij}$, that is, a Kronecker delta. The quantities $C_{ij}$ are the elements of the confusion matrix, $\mathbf{C}$, which has dimension $n\times n$.
The primary drawback of imperfect classification arises when studying the characteristics of galaxy populations using predicted classes, a method extensively used in the literature. Without any knowledge of $\mathbf{C}$, results from such studies can be misleading. Let us suppose we are interested in determining the distribution of a particular galaxy property $x$ (e.g. colour, specific star formation rate, size, etc.) for galaxies of a given stellar mass, or, alternatively, absolute magnitude, $M$. The observed distributions of the property $x$ are obtained by counting galaxies of a predicted class in $m$ bins of $x$ at a fixed $M$. The fraction of galaxies of the $i-$th predicted class with $x_{k}-\Delta x_{k}/2\leq x\leq x_{k}+\Delta x_{k}/2$ and normalised to have unity area when integrating over the whole range of $x$ is:
@@
where the intrinsic (desired) quantities are the $F_{jk}(M)$. The normalisation condition is:
@@
If we have an estimation of $\mathbf{C}$, for instance, by using numerical simulations, we can use Eq. 2 to solve for $F_{jk}(M)$. By arranging the quantities $f_{ik}(M)$ in a $n\times m$ matrix, $\mathbf{f}$, with rows $i$ (classes) and columns $k$ (bins in $x$), and doing the same for $F_{jk}(M)$ in $\mathbf{F}$, now Eq. 2 reads:
@@
It is important to remark that, so far, we are assuming the confusion matrix does not depend on mass (or absolute magnitude). In Sect. 3.3 below, we explore whether a stellar mass dependence of the confusion matrix (i.e. $\mathbf{C}\equiv\mathbf{C}(M)$) provides better results.
@@ 3.1 The sample of simulated clusters and galaxies
We use the same sample of clusters and galaxies used by de los Rios et al. (2021), to train and test the code roger. The sample of cluster includes 34 massive systems, $M_{200}\geq 10^{15}h^{-1}M_{\odot}$, from the MultiDark Planck 2 simulation (Klypin et al. 2016). These clusters were selected to ensure that they do not have a companion halo more massive than 10 per cent of $M_{200}$ within $5\times R_{200}$. The mass $M_{200}$ is estimated by considering an homogeneous mass distribution enclosed by a characteristic radius of a dark matter halo, $R_{200}$, at which the density is equal to $200$ times the critical density of the Universe at the redshift of the system.
@@
    Cluster members (CL): They may have crossed $R_{200}$ several times in the past and are now found orbiting around the cluster centre, with the majority of them within $R_{200}$ of the centre.
@@
    Backsplash galaxies (BS): Galaxies that have crossed twice $R_{200}$. The first time diving in, and the second time on their way out of the cluster, where they are now. Most of them will fall back into the cluster in the future.
@@
    Recent infallers (RIN): Galaxies that have crossed $R_{200}$ only once as they dived into the cluster in the past 2 Gyr. Many of them may be BS in the future.
@@
    Infalling galaxies (IN): Galaxies that have always been at distances greater than $R_{200}$ from the cluster centre, and have negative radial velocity relative to the cluster. These galaxies are falling to the cluster.
@@
    Interlopers (ITL): Galaxies that have never got closer than $R_{200}$ from the cluster centre, but unlike IN, they have positive radial velocity relative to the cluster, i.e. they are receding away from the cluster at $z=0$. We consider these galaxies as objects that will not fall into the cluster. They are unrelated to the cluster but can be confused with classes 1–4 above in the PPSD, i.e., they are contamination.
Of the total of 30,289 galaxies, we use half of them to re-train roger, the training set, and the other half is the test set, which is the subsample of galaxies we use to compute the confusion matrix $\mathbf{C}$. In de los Rios et al. (2021), we used $\sim 80\%$ of the complete sample as training set, and the remaining as the test set. Here, we are interested in having a larger test set; thus, for the sake of consistency, we performed the re-training of roger accordingly. It is noteworthy that we are using a python version of roger  which we dub pyroger that is publicly available at roger’s website.
For each galaxy in the test set, we used pyroger to compute its probability of being of each class, $p_{i}$ where $i=1,\ldots,5$, and the correspondence between index and class is as listed above: $1\leftrightarrow$ CL, $2\leftrightarrow$ BS, $3\leftrightarrow$ RIN, $4\leftrightarrow$ IN, and $5\leftrightarrow$ ITL. roger computes these probabilities using three different techniques: K-Nearest Neighbours, Support Vector Machine, and Random Forest. We refer the reader to de los Rios et al. (2021) for details. For the purposes of this work, we use the probabilities pyroger computes using the K-Nearest Neighbours technique.
We use the classification scheme adopted by Coenda et al. (2022): the predicted class is defined by the highest value of $p_{i}$, and also this value has to be higher than a class-dependent threshold. These thresholds are: 0.4, 0.48, 0.37, 0.54, and 0.15, for CL, BS, RIN, IN, and ITL, respectively. This scheme is an appropriate balance between sensitivity and precision.
@@
Figure 2: Colour distributions of galaxies. Each class is shown in a different row, as quoted at the right hand side of the figure. Each column considers a particular range of galaxy stellar mass as quoted on top of the figure. Gray shaded histograms are the real colour distributions, i.e., they correspond to the intrinsic classes. Green lines are the colour distributions of the predicted classes. Violet lines are the colour distributions recovered by our method. We quote within each panel the values of the corresponding square residuals, $S^{\rm pred}$ and $S^{\rm rec}$, given by Eqs. 6 and 7, respectively.
Figure 3: Sum of squares residuals as a function of galaxy stellar mass. Each panel considers a different galaxy class as quoted inside. Residuals between the real and the predicted distributions are shown in green solid lines, i.e., $S^{\rm pred}$ in eq. 6. Residuals between the real and the recovered distributions are shown in violet solid lines, i.e., $S^{\rm rec}$ in eq. 7. We show in solid orange lines the residuals between the real and the recovered distributions when we use a stellar mass dependent confusion matrix. Note that the $y-$axis scale varies from panel to panel.
@@ 3.2 Applying the method
As a first step, we construct the confusion matrix using the test set, and compute $C_{ij}$ as the fraction of galaxies of predicted class $i$ that are intrinsically of class $j$. For a better visualisation, instead of simply providing the numbers, we show the confusion matrix in Fig. 1. We can see in Fig. 1 how the predicted classes are composed in terms of the real classes. Predicted ITL are the simplest case; they are basically real ITL with a mild contamination ($13\%$) of IN. Predicted IN are roughly real IN with two sources of contamination: $30\%$ ITL and 9% BS. As for predicted RIN, they are 52% real RIN, with an unsurprising main source of contamination from CL (23%), and lower ($\leq 10\%$) percentages from the other classes. Predicted BS are $53\%$ real BS and their main contamination comes from misclassified IN ($32\%$). Finally, only $44\%$ predicted CL are real CL; their most significant contamination is from RIN ($26\%$), and in second place BS and IN contributing $\sim 10\%$ each. We refer the reader to de los Rios et al. (2021) for a thorough discussion on how the contamination between the five classes works. The three most contaminated predicted classes, CL, BS, and IN, are the most interesting ones for studying the effects of environments in the evolution of galaxies.
In Fig. 2, we present the distributions of $x=u-r$ colour for the intrinsic classes, the predicted classes, and the reconstructed distributions obtained through the inversion of the confusion matrix, according to the adopted classification scheme. We split galaxies into five stellar mass bins in our analysis. To quantify whether our method provides better estimations of the intrinsic colour distribution, we compute the sum of squared residuals for each class $i$. This is done by comparing the intrinsic distributions $F_{ik}(M)$ (where $k=1,\ldots,m$ denote the bins in colour) with the colour distributions of both the predicted classes, $f_{ik}(M)$, and the distributions recovered by our method, $F_{ik}^{\rm rec}(M)$, i.e.
@@
Note that we compute the square residuals and not $\chi^{2}$ since there are bins where $F_{ik}(M)=0$. These quantities are quoted inside each panel in Fig. 2. For a better visualisation and comparison, we show in Fig. 3 the sum of square residuals as a function of stellar mass. Note that $y-$axis scale varies from panel to panel in Fig. 3.
A joint analysis of Figs. 2 and 3 gives two general conclusions: on the one hand, our method improves the determination of the colour distribution of the three predicted classes that are more affected by misclassifications, namely, CL, BS, and RIN; and on the other hand, it does not improve the results for IN and ITL, which are the two predicted classes best classified by our adopted criteria. These two main results are promising for this type of study; however, there is more to consider beyond just two levels of contamination, namely high contamination for CL, BS and RIN and low contamination for IN and ITL. Regarding their intrinsic colour distributions, the IN and ITL classes are the most similar, while the other three classes show greater variations from one to another. This is the consequence of very different evolutionary paths. Delving into the specifics, there are several interesting details in Figs. 2 and 3, which we summarise below.
@@
    Cluster members: Our method improves greatly the determination of the colour distributions in all five stellar mass bins. The predicted class CL exhibits a contamination from blue galaxies, which is particularly strong at the lower mass bins. These galaxies are mostly RIN, and in a second place BS, incorrectly classified as CL. Distributions recovered by our method eliminate most of this contamination. We consistently find $S_{\rm CL}^{\rm rec}(M)<S_{\rm CL}^{\rm pred}(M)$ for all mass bins.
@@
    Backsplash galaxies: As in the previous case, there are clear improvements for BS galaxies too, consisting in the removal of the contamination from blue, and mainly IN, galaxies. For all mass bins we find $S_{\rm BS}^{\rm rec}(M)<S_{\rm BS}^{\rm pred}(M)$. Backsplash galaxies are, on average, redder when contamination effects are removed by our method.
@@
    Recent infallers: With the exception of the highest mass bin, we find that the method produces better results, i.e., $S_{\rm RIN}^{\rm rec}(M)<S_{\rm RIN}^{\rm pred}(M)$. The improvement is most notably seen in the removal of the contamination from misclassified red CL galaxies with stellar mass in the range $9.7\leq\log(M/h^{-1}M_{\odot})\leq 10.5$ (i.e. third and fourth columns in Fig. 2). Our findings for CL above and for RIN here suggest that almost every blue galaxy found in a cluster should have dived in there recently. This could have implications on the characterisation of satellites in clusters according to their dynamical state (e.g. Aldás et al. 2023, 2024).
@@
    Infalling galaxies: In this case the method gives results comparable to the raw predictions. We can observe no significant differences between the intrinsic, the predicted, and the recovered distributions in Fig. 3. The values of the sum of square residuals are $S_{\rm IN}^{\rm rec}(M)>S_{\rm IN}^{\rm pred}(M)$ for three alternate mass bins, and otherwise in the remaining two. Note that both, $S_{\rm IN}^{\rm rec}(M)$ and $S_{\rm BS}^{\rm pred}(M)$, are very similar in all cases and take values $\sim 1-2\times 10^{-2}$, which is, on average, smaller than the corresponding values we find for CL, BS, and RIN.

### v2

# Improving the accuracy of observable distributions for galaxies classified in the projected phase space diagram

@@ Abstract
Aims. We propose a method for statistically correcting the determination of galaxy properties’ distributions that accounts for the contamination caused by misclassified galaxies from other classes.
Methods. Using a sample of massive clusters and the galaxies in their surroundings taken from the MultiDark Planck 2 simulation combined with the semi-analytic model of galaxy formation sag, we computed the confusion matrix associated with a classification scheme in the PPSD. Based on positions in the PPSD, galaxies are classified as cluster members, backsplash galaxies, recent infallers, infalling galaxies, or interlopers. This classification is determined using probabilities calculated by the code roger  along with a threshold criterion. By inverting the confusion matrix, we are able to get better determinations of distributions of galaxy properties, such as colour.
Results. Compared to a direct estimation based solely on the predicted galaxy classes, our method provides better estimates of the mass-dependent colour distribution for the galaxy classes most affected by misclassification: cluster members, backsplash galaxies, and recent infallers. We applied the method to a sample of observed X-ray clusters and galaxies.
Conclusions. Our method can be applied to any classification of galaxies in the PPSD, and to any other galaxy property besides colour, provided an estimation of the confusion matrix is available. Blue, low-mass galaxies in clusters are almost exclusively recent infaller galaxies that have not yet been quenched by the environmental action of the cluster. Backsplash galaxies are on average redder than expected.
@@ 1 Introduction
In recent years, the projected phase space diagram (PPSD) has increasingly been used as a tool for classifying galaxies according to their kinematic behaviours and dynamical state in and around systems of galaxies, such as clusters and groups. The PPSD is a 2D space that combines the projected cluster-centric (or group-centric) distance with the line-of-sight velocity relative to the cluster (or group). Several authors have utilised the PPSD to study the evolution of galaxies. For example, Mahajan et al. (2011) studied the decrease in the star formation of backsplash galaxies (BSs). Muzzin et al. (2014) focused on the quenching of galaxies at $z\sim 1$. Muriel & Coenda (2014) investigated the properties of high- and low-velocity galaxies in the outskirts of galaxy clusters. Hernández-Fernández et al. (2014) and Jaffé et al. (2015) studied the gas fraction and ram pressure of galaxies in clusters. Oman & Hudson (2016) studied the relationship between the star formation rates of a sample of observed galaxies and their likely orbital histories. Smith et al. (2019) analysed the stellar mass growth histories of galaxies as a function of their infall time. Martínez et al. (2023) studied the quenching of galaxies in a sample of X-ray clusters. Aguerri et al. (2023) focused on the properties of barred galaxies and their environments. Sampaio et al. (2024) examined the evolution of blue cloud, green valley, and red-sequence fractions as a function of time since infall. Muñoz Rodríguez et al. (2024) studied the impact on black hole growth as a function of the environment.
@@
The aforementioned studies classify galaxies in systems and their outskirts following methods such as those proposed by Rhee et al. (2017), Pasquali et al. (2019), or de los Rios et al. (2021). Based on cosmological hydrodynamic N-body simulations of groups and clusters (Choi & Yi 2017), Rhee et al. (2017) categorised galaxies within the PPSD based on the time since their infall into the system. This categorisation included galaxies identified as first, recent, intermediate, and ancient infallers. They delineated specific regions within the PPSD where each of these types of galaxies are preferentially located. Pasquali et al. (2019) present an alternative approach that uses the same cosmological simulations (Choi & Yi 2017) of groups and clusters employed by Rhee et al. (2017). They established eight zones of constant mean infall time to investigate environmental influences on satellite galaxies. These zones are defined by analytical curves that characterise the timing of galaxy infall into the system.
@@
Recently, de los Rios et al. (2021) presented the code roger [^1] https://github.com/Martindelosrios/ROGER, which employs three different machine learning techniques to classify galaxies based on their projected phase-space positions within and around clusters. They utilised a dataset comprising massive and isolated galaxy clusters from the MultiDark Planck 2 simulation (Klypin et al. 2016). This code establishes a connection between the 2D phase space position of galaxies and their 3D orbital classifications. roger calculates the probabilities of galaxies belonging to any of the five orbital classes: cluster members (CLs), BSs, recent infallers (RINs), infallers (INs), and interlopers (ITLs).
@@
Regardless of the chosen method for classifying galaxies in the PPSD, there will inevitably be some degree of contamination among the classes, particularly in regions closer to the centres of clusters. According to Coenda et al. (2022), misclassifications have significant implications for interpreting the properties of galaxies in the 2D classes. Therefore, the degree of contamination significantly impacts the conclusions that can be drawn. They emphasise the importance of distinguishing between red and blue galaxy populations to achieve a more precise analysis of observational data. Furthermore, they argue that the 2D analysis reliably provides results only for RINs, INs, and ITLs within the blue population.
@@
To improve determinations of distributions of galaxy observables, such as colours or the specific star formation rate, for galaxies that have been classified using their position in the PPSD, we propose a simple method that reconstructs the intrinsic distribution of galaxy properties. The method relies on a statistic estimation of how well the classification is performed. This paper is organised as follows: In Sect. 2 we present our method; in Sect. 3 we test the method on a sample of simulated clusters and galaxies; in Sect. 4 we use the method to reconstruct the distributions of colours of galaxies of different classes in and around a sample of X-ray clusters of galaxies; finally, we present our conclusions and mention possible further uses of the method in Sect. 5.
@@ 2 Method
If we consider $n$ classes of galaxies according to the particular classification scheme based on the PPSD we are using, the intrinsic class of each galaxy cannot be determined based solely on its position on it. At best, we can assign a particular class to a galaxy based on its position in the PPSD using our chosen methodology. In our case, this involves utilising the output probabilities of roger combined with a numerical criterion on their values, as we explain in Sect. 3. This way, each galaxy is assigned a predicted class. Due to uncertainties inherent to the classification procedure, the number of galaxies in the $i-$th predicted class ($i=1,\ldots,n$), $N^{\mathrm{P}}_{i}$, is built up from the contribution of galaxies, in principle from all intrinsic classes, that were classified as class $i$:
@@
where $N^{\mathrm{I}}_{j}$ is the number of galaxies in our sample that belong to the $j-$th intrinsic class, and $C_{ij}$ is the fraction of galaxies of the $j-$th intrinsic class that were classified as being part of the $i-$th predicted class. The ideal case is $C_{ij}=\delta_{ij}$, that is, a Kronecker delta. The quantities $C_{ij}$ are the elements of the confusion matrix, $\mathbf{C}$, which has dimensions $n\times n$.
The primary drawback of imperfect classification arises when studying the characteristics of galaxy populations using predicted classes, a method extensively used in the literature. Without any knowledge of $\mathbf{C}$, results from such studies can be misleading. If we want to determine the distribution of a particular galaxy property, $x$ (e.g. colour, specific star formation rate, size, etc) for galaxies of a given stellar mass or, alternatively, absolute magnitude ($M$), the observed distributions of the property, $x$, are obtained by counting galaxies of a predicted class in $m$ bins of $x$ at a fixed $M$. The fraction of galaxies of the $i-$th predicted class with $x_{k}-\Delta x_{k}/2\leq x\leq x_{k}+\Delta x_{k}/2$ and normalised to have unity area when integrating over the whole range of $x$ is
@@
where the intrinsic (desired) quantities are the $F_{jk}(M)$. The normalisation condition is
@@
If we have an estimation of $\mathbf{C}$, for instance, by using numerical simulations, we can use Eq. 2 to solve for $F_{jk}(M)$. By arranging the quantities $f_{ik}(M)$ in a $n\times m$ matrix, $\mathbf{f}$, with rows $i$ (classes) and columns $k$ (bins in $x$), and doing the same for $F_{jk}(M)$ in $\mathbf{F}$, now Eq. 2 reads
@@
It is important to remark that so far we have assumed that the confusion matrix does not depend on mass (or absolute magnitude). In Sect. 3.3 we explore whether a stellar mass dependence of the confusion matrix (i.e. $\mathbf{C}\equiv\mathbf{C}(M)$) provides better results.
@@ 3.1 The sample of simulated clusters and galaxies
We used the same sample of clusters and galaxies used by de los Rios et al. (2021) to train and test the code roger. The sample of cluster includes 34 massive systems, $M_{200}\geq 10^{15}h^{-1}M_{\odot}$, from the MultiDark Planck 2 simulation (Klypin et al. 2016). These clusters were selected to ensure that they do not have a companion halo more massive than 10 per cent of $M_{200}$ within $5\times R_{200}$. The mass $M_{200}$ is estimated by considering an homogeneous mass distribution enclosed by a characteristic radius of a dark matter halo, $R_{200}$, at which the density is equal to $200$ times the critical density of the Universe at the redshift of the system.
@@
    Cluster members: They may have crossed $R_{200}$ several times in the past and are now found orbiting around the cluster centre, with the majority of them within $R_{200}$ of the centre.
@@
    Backsplash galaxies: Galaxies that have crossed twice $R_{200}$. The first time diving in, and the second time on their way out of the cluster, where they are now. Most of them will fall back into the cluster in the future.
@@
    Recent infallers: Galaxies that have crossed $R_{200}$ only once as they dived into the cluster in the past 2 Gyr. Many of them may be BS in the future.
@@
    Infalling galaxies: Galaxies that have always been at distances greater than $R_{200}$ from the cluster centre, and have negative radial velocity relative to the cluster. These galaxies are falling to the cluster.
@@
    Interlopers: Galaxies that have never gotten closer than $R_{200}$ from the cluster centre but, unlike INs, have positive radial velocities relative to the cluster, that is, they are receding away from the cluster at $z=0$. We consider these galaxies as objects that will not fall into the cluster. They are unrelated to the cluster but can be confused with classes 1–4 above in the PPSD (i.e. they are contamination).
Of the total of 30,289 galaxies, we used half of them to re-train roger (the training set), and the other half is the test set, which is the subsample of galaxies we used to compute the confusion matrix $\mathbf{C}$. In de los Rios et al. (2021), we used $\sim 80\%$ of the complete sample as training set, and the remaining as the test set. Here, we are interested in having a larger test set; thus, for the sake of consistency, we re-trained roger accordingly. It is noteworthy that we used a python version of roger, which we dub pyroger, that is publicly available at roger’s website.
For each galaxy in the test set, we used pyroger to compute its probability of being a member of each class, $p_{i}$ where $i=1,\ldots,5$, and the correspondence between index and class is as listed above: $1\leftrightarrow$ CL, $2\leftrightarrow$ BS, $3\leftrightarrow$ RIN, $4\leftrightarrow$ IN, and $5\leftrightarrow$ ITL. roger computes these probabilities using three different techniques: K-nearest neighbours, support vector machine, and random forest. We refer the reader to de los Rios et al. (2021) for details. For the purposes of this work, we used the probabilities that pyroger computes using the K-nearest neighbours technique.
We used the classification scheme adopted by Coenda et al. (2022): the predicted class is defined by the highest value of $p_{i}$, and this value has to be higher than a class-dependent threshold. These thresholds are 0.4, 0.48, 0.37, 0.54, and 0.15, for CLs, BSs, RINs, INs, and ITLs, respectively. This scheme is an appropriate balance between sensitivity and precision.
@@
Figure 2: Colour distributions of galaxies. Each class is shown in a different row, as noted to the right of the panels. Each column considers a particular range of galaxy stellar mass, as noted at the top. Grey shaded histograms are the real colour distributions, i.e. they correspond to the intrinsic classes. Green lines are the colour distributions of the predicted classes. Violet lines are the colour distributions recovered by our method. We quote within each panel the values of the corresponding square residuals, $S^{\rm pred}$ and $S^{\rm rec}$, given by Eqs. 6 and 7, respectively.
Figure 3: Sum of square residuals as a function of galaxy stellar mass. Each panel considers a different galaxy class. Residuals between the real and the predicted distributions are shown with green lines, i.e. $S^{\rm pred}$ in Eq. 6. Residuals between the real and the recovered distributions are shown with violet lines, i.e. $S^{\rm rec}$ in Eq. 7. The residuals between the real and the recovered distributions when we use a stellar mass-dependent confusion matrix are shown in orange. Note that the $y-$axis scale varies from panel to panel.
@@ 3.2 Applying the method
As a first step, we constructed the confusion matrix using the test set and computed $C_{ij}$, the fraction of galaxies of predicted class $i$ that are intrinsically of class $j$. For a better visualisation, instead of simply providing the numbers, we show the confusion matrix in Fig. 1. We can see in Fig. 1 how the predicted classes are composed in terms of the real classes. Predicted ITLs are the simplest case; they are basically real ITLs with a mild contamination ($13\%$) of INs. Predicted INs are roughly real INs with two sources of contamination: $30\%$ ITLs and 9% BSs. As for the predicted RINs, 52% are real RINs, with an unsurprising main source of contamination from CLs (23%) and lower ($\leq 10\%$) percentages from the other classes. Of the predicted BSs, $53\%$ are real BSs, and their main contamination comes from misclassified INs ($32\%$). Finally, only $44\%$ of the predicted CLs are real CLs; their most significant contamination is from RINs ($26\%$), and in second place BSs and INs contributing $\sim 10\%$ each. We refer the reader to de los Rios et al. (2021) for a thorough discussion on how the contamination between the five classes works. The three most contaminated predicted classes, CLs, BSs, and INs, are the most interesting ones for studying the effects of environments in the evolution of galaxies.
In Fig. 2 we present the distributions of $x=u-r$ colour for the intrinsic classes, the predicted classes, and the reconstructed distributions obtained through the inversion of the confusion matrix, according to the adopted classification scheme. We split galaxies into five stellar mass bins in our analysis. To quantify whether our method provides better estimations of the intrinsic colour distribution, we computed the sum of squared residuals for each class $i$. This was done by comparing the intrinsic distributions ($F_{ik}(M)$, where $k=1,\ldots,m$ denote the bins in colour) with the colour distributions of both the predicted classes ($f_{ik}(M)$) and the distributions recovered by our method, $F_{ik}^{\rm rec}(M)$:
@@
We note that we computed the square residuals and not $\chi^{2}$ since there are bins where $F_{ik}(M)=0$. These quantities are quoted inside each panel in Fig. 2. For a better visualisation and comparison, we show in Fig. 3 the sum of square residuals as a function of stellar mass. We note that $y-$axis scale varies from panel to panel in Fig. 3.
A joint analysis of Figs. 2 and 3 gives two general conclusions: on the one hand, our method improves the determination of the colour distribution of the three predicted classes that are more affected by misclassifications, namely, CLs, BSs, and RINs; and on the other hand, it does not improve the results for INs and ITLs, which are the two predicted classes best classified by our adopted criteria. These two main results are promising for this type of study; however, there is more to consider beyond just two levels of contamination, namely high contamination for CLs, BSs, and RINs and low contamination for INs and ITLs. Regarding their intrinsic colour distributions, the IN and ITL classes are the most similar, while the other three classes show greater variations from one to another. This is the consequence of very different evolutionary paths. Delving into the specifics, there are several interesting details in Figs. 2 and 3, which we summarise below.
@@
    Cluster members: Our method improves greatly the determination of the colour distributions in all five stellar mass bins. The predicted CL class exhibits a contamination from blue galaxies, which is particularly strong in the lower-mass bins. These galaxies are mostly RINs, and in a second place BSs, incorrectly classified as CLs. Distributions recovered by our method eliminate most of this contamination. We consistently find $S_{\rm CL}^{\rm rec}(M)<S_{\rm CL}^{\rm pred}(M)$ for all mass bins.
@@
    Backsplash galaxies: As in the previous case, there are clear improvements for BS galaxies too, consisting in the removal of the contamination from blue, and mainly INs, galaxies. For all mass bins we find $S_{\rm BS}^{\rm rec}(M)<S_{\rm BS}^{\rm pred}(M)$. Backsplash galaxies are, on average, redder when contamination effects are removed by our method.
@@
    Recent infallers: With the exception of the highest-mass bin, we find that the method produces better results (i.e. $S_{\rm RIN}^{\rm rec}(M)<S_{\rm RIN}^{\rm pred}(M))$. The improvement is most notably seen in the removal of the contamination from misclassified red CL galaxies with stellar mass in the range $9.7\leq\log(M/h^{-1}M_{\odot})\leq 10.5$ (see the third and fourth columns in Fig. 2). Our findings for CLs above and for RINs here suggest that almost every blue galaxy found in a cluster should have dived in there recently. This could have implications on the characterisation of satellites in clusters according to their dynamical state (e.g. Aldás et al. 2023, 2024).
@@
    Infalling galaxies: In this case the method gives results comparable to the raw predictions. We can observe no significant differences between the intrinsic, the predicted, and the recovered distributions in Fig. 3. The values of the sum of square residuals are $S_{\rm IN}^{\rm rec}(M)>S_{\rm IN}^{\rm pred}(M)$ for three alternate mass bins, and otherwise in the remaining two. We note that $S_{\rm IN}^{\rm rec}(M)$ and $S_{\rm BS}^{\rm pred}(M)$ are very similar in all cases and take values $\sim 1-2\times 10^{-2}$, which is, on average, smaller than the corresponding values we find for CLs, BSs, and RINs.